Fermi National Laboratory
 

Operation Summary for February 4, 2011 - February 7, 2011

 

The Integrated Luminosity for the period from 1/31/11 to 2/7/11 was 43.64 Pb-1 (inverse picobarns). NuMI reported receiving 5.14E18 protons on target during this same period.

Friday Feb 4

The day shift began with Operations monitoring the Tevatron (TEV) store 8473 (49.1E30) and the Recycler stash (118.5E10), with the Antiproton Source (Pbar) stacking (27.4E10), with the Linac, Booster, and Main Injector (MI) operating normally, with the NuMI, MiniBooNE, and Meson FTBF T-978 (CALICE) experiments taking beam, and with the Collider Detector at Fermilab (CDF) and D-Zero (D0) using the colliding beams.

During the day shift, Operations reported that store 8473 aborted at 12:42 PM due to a bad relay in the TEV sector D3 VCB cabinet. EE Support personnel also reported finding a bad relay in the A2 VCB cabinet; they spent the rest of the shift checking relays around the ring. EE Support personnel changed the bad NuMI VCB out and Operations restored beam at 1:12 PM. Operations disabled beam to T-978 for an access at 1:29 PM. Operators prepared TEV sectors A-E for access from 2:38 PM through to the end of shift. Operations reported a short downtimes due to JASMIN experimenters retrieving a sample from Pbar, to a Booster RF station trip (BRF13), and due to a NuMI beam permit trip.

During the evening shift, Operations reported that Mechanical technicians had accessed the TEV, and that EE Support personnel were conducting VCB inspections; the VCB inspections halted beam to Pbar and NuMI for two hours. Mechanical techs came out of TEV A-E access at 5:22 PM. Operators began the TEV turn on at 6:40 PM. Operations put the TEV into HEP shot setup at 10:28 PM and established store 8477 at 11:25 PM with an average initial luminosity (AIL) of 339.25E30. Operations reported some Meson Cryo Central power issues at the end of shift, that BRF14 had started tripping on repeated gap envelope faults, and that MIRF10 had faulted off requiring operators to investigate locally.

Saturday Feb 5

The midnight shift began with Operations monitoring store 8477 (270.89E30) and the stash (80.7E10) and the Pbar stack (4.2E10), and with NuMI and MiniBooNE taking beam.

During the midnight shift, Operations reported that FESS personnel continued to investigate the Meson Cryo Central power outage. Operations reported that MIRF10 tripped off and would not turn on; operators bypassed the station and adjusted the RF system to run without the station. A FESS electrician reported finding a 480 V 2 kA breaker tripped. He reset the breaker and CHL personnel powered up the Meson Cryo Central systems. Operations established beam to T-978. MIRF14 tripped off at 7:02 AM and would not run stably when reset. MIRF techs were en route to investigate by the end of shift.

During the day shift, Operations terminated store 8477 at 3:19 PM.

During the evening shift, Operations established store 8478 at 4:30 PM with an AIL of 318E30. Operators rebooted the A1 BPMs. RF techs reported that MIRF10 was broken in the tunnel. Switchyard (SY) personnel conducted splitting studies. Operations disabled beam to T-978 at 10 PM.

Sunday Feb 6

The midnight shift began with Operations monitoring store 8478 and the stash, with Pbar stacking, and with NuMI and MiniBooNE taking beam.

During the midnight shift, at 3:12 AM, cryo system personnel at CHL reported that cold box #2 had tripped off. Operations terminated store 8478 when CHL reported that they could not maintain liquid helium in the satellite refrigerator buildings. Operators dumped the ramp and de-energized the low beta magnets. Operations established beam to T-978 at 4:52 AM. CHL reported that they had the cold box repaired and had started recovery by the end of shift.

During the day shift, Operations was waiting for CHL recovery. Operators turned the TEV on to a QBS ramp at 3:30 PM and left it at 90 GeV pending a High Energy ramp permit. The stash hit its Recycler limit early in shift. Operations turned off Pbar for 2 hours due to beam detector testing. Operations reported that a Booster kicker (MKS05) had sparked and tripped off and had also caused BRF17 to trip off. Booster Personnel came in and made repairs. Operations had Booster beam by end of shift.

During the evening shift, Booster personnel re-terminated the MKS05 PFN cable and had it back online by 9:50 PM. Operations disabled beam to T-978; this ended their run. CHL personnel deemed that the A3 cryo system was cold at 10:51 PM. Operators began the TEV turn-around.

Monday Feb 7

The midnight shift began with Operations monitoring the stash, with Pbar stacking, and with NuMI and MiniBooNE taking beam.

During the midnight shift, Operations put the TEV into shot setup at 1:23 AM and established store 8482 at 2:39 AM with an AIL of 356.12e30. Operations prepared the SY F1, F2, and F3 manholes for access. Booster RF techs brought BRF17 online by at the end of the midnight shift.

The Plans for Monday

The plans for today are to stack, stash, and store.

 

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 02/07/2011    email Fermilab
Security, Privacy, LegalFermi National Accelerator Laboratory